# Fairness in Serving Large Language Models

### v1

@@ Abstract
High-demand LLM inference services (e.g., ChatGPT and BARD) support a wide range of requests from short chat conversations to long document reading. To ensure that all client requests are processed fairly, most major LLM inference services have request rate limits, to ensure that no client can dominate the request queue. However, this rudimentary notion of fairness also results in under-utilization of the resources and poor client experience when there is spare capacity. While there is a rich literature on fair scheduling, serving LLMs presents new challenges due to their unpredictable request lengths and their unique batching characteristics on parallel accelerators.
This paper introduces the definition of LLM serving fairness based on a cost function that accounts for the number of input and output tokens processed. To achieve fairness in serving, we propose a novel scheduling algorithm, the Virtual Token Counter (VTC), a fair scheduler based on the continuous batching mechanism. We prove a 2×2$\times$2 × tight upper bound on the service difference between two backlogged clients, adhering to the requirement of work-conserving. Through extensive experiments, we demonstrate the superior performance of VTC in ensuring fairness, especially in contrast to other baseline methods, which exhibit shortcomings under various conditions.
@@ 1 Introduction
Today’s LLM serving systems [20, 16] typically use First-Come-First-Serve (FCFS) to schedule incoming requests. While simple, this scheduling discipline has several drawbacks. One such drawback is the lack of isolation: a client sending a disproportionate number of requests can negatively impact the service of all the other clients sharing the same server (i.e., slow down their requests or even cause timeouts) even when they send very little traffic. One solution to address this problem is to limit the incoming load of each client. Many of the existing LLM services do this today by imposing a request-per-minute (RPM) limit [33] for each client.
@@
[…]
@@
Despite these similarities, we cannot directly use the algorithms developed for networking and operating systems, as LLM serving has several unique characteristics. First, the request output lengths are unknown in advance. In contrast, in networking, the packet lengths are known before the packet is scheduled. Second, the cost of each token can vary. For instance, the cost of processing an input (prompt) token is typically lower than that of an output token, because input tokens processing is parallelizable. In contrast, the cost of sending a bit or the cost of a CPU time slice are the same irrespective of the workload. Third, the effective capacity of an LLM server (i.e., processing rate expressed in token/sec when the request queue is non-empty) can vary over time. For example, longer input sequences take more memory. This limits the number of batched parallel requests during generation, leading to GPU under-utilization and a lower processing rate. In contrast, the network or CPU capacity is assumed to be fixed.
In this paper, we discuss the factors that need to be considered when defining fairness in the context of LLM serving. We show how different definitions can be incorporated into a configurable cost function (Section 3). We then present a fair scheduling algorithm called Virtual Token Counter (VTC) that can be easily adapted for different cost functions. At a high level, VTC tracks the services received for each client and will prioritize the ones with the least services received, with a counter lift each time a client is new to the queue. It updates the counters at a token-level granularity on the fly, which addresses the unknown length issue. VTC integrates seamlessly with current LLM serving batching techniques (Section 2.1), and its scheduling mechanism does not depend on the server’s capacity, overcoming the problem of the dynamically fluctuating server capacity. We also provide theoretical bounds of fairness for VTC in Section˜4.1. The system architecture of VTC implementation in illustrated in Figure 1.
@@
  • This is the first work to discuss the fair serving of Large Language Models. We identify its unique challenges and give the definition of LLM serving fairness (Section˜3).
  • We propose a simple yet effective fair-serving algorithm called VTC. We provide rigorous proofs for VTC on fairness guarantee, which gives fairness bound within 2×2$\times$2 × of the optimal bound (Section˜4).
  • We conduct in-depth evaluations on our proposed algorithm VTC. Results confirm that our proposed algorithms are fair and work-conserving, while other alternatives can fail in different settings (Section˜5).
@@ 2.1 Large Language Models Serving
The generation procedure consists of two stages: the initial prefilling stage, and the decoding stage [35]. Mathematically, $x$ is a sequence of tokens $(x_{1},x_{2},...,x_{n})$. In the prefilling stage, the LLM computes the probability of the first new tokens: $P(x_{n+1}|x_{1},...,x_{n})$. In the decoding stage, the system autoregressively generates a new token. At time $t$ ($t\geq 1$), the process is written as:
$P(x_{n+t+1}|x_{1},...,x_{n+t})$
@@
Figure 1: Serving architecture with Virtual Token Counter (VTC), illustrated with two clients. VTC maintains a queue of requests and keeps track of tokens served for each client. In each iteration of the LLM execution engine, some tokens from some clients are generated. The counters of these clients are correspondingly updated. When the condition of adding new requests is satisfied (e.g. memory is released when some other requests finish), VTC will be invoked to choose the requests to be added. VTC achieves fairness by prioritizing clients with the lowest counter (Section 4.1).
@@
Naively, the execution stream can choose to execute requests one by one. However, this is highly GPU inefficient due to various natures of the LLM generation procedure. For instance, the decoding steps must be carried out sequentially where the arithmetic intensity is relatively low in a single step. Contemporary serving systems usually perform batching that executes multiple requests concurrently to maximize the system throughput. The most widely used approach in LLM serving is continuous batching [46]. Algorithm 1 shows the pseudocode for continuous batching. [^2] For a simple presentation, we consider the continuous batching from TGI [17] rather than the original proposed iteration-level scheduling in Orca [46]. The monitoring stream enqueues requests to a waiting queue. The execution stream performs a check on whether there are finished requests at the end of each decoding step. If there are, the system removes these requests and adds new requests from the queue.
@@
We adopt a continuous batching scheme where a request only leaves the batch when it generates an EOS token or reaches the pre-defined maximum number of generated tokens (i.e., no preemption). This paper focuses on how to integrate fair scheduling with continuous batching, and we leave an investigation on preemption as an orthogonal future work.
@@ 2.2 Existing Fairness Approaches
Virtually all fair queueing algorithms maintain a system virtual time, $v(t)$, which intuitively measures the service received by a continuously backlogged flow in terms of bits forwarded. Each packet, $p$ is associated two tags: Start tag $S(p)$ and a Finish tag $F(p)=S(p)+l_{p}$. The Start tag (a.k.a. packet’s virtual starting time) is computed based on both the system virtual time and the Finish tag (a.k.a. packet’s virtual finishing time). These algorithms schedule packets in the ascending order of either the Finish or Start tags.
@@
There exists a distinct strand of research [4, 2, 43] focusing on the fair scheduling of preemptible tasks (e.g., CPU scheduling). However, our setting for LLM serving aligns more closely with the packet scheduling problem, and introduces new challenges such as unpredictable output length, cost disparity between input and output tokens, and variable token-rate servers.
@@ 2.3 Challenges
Additionally, in LLM serving, the server has variable token-rate capacity, although the memory allocated for a batch is constant. Firstly, even if the request queue is not empty, we are not guaranteed that each batch is full. This is because we need to preserve spaces for future generated tokens, and also because the tokens added to the batch are not at the token but the request granularity. Secondly, the number of tokens processed highly depends on the requests’ arrival patterns because of the continuous batching mechanism (Section 2.1). Furthermore, the capacity depends on the mix between input and output tokens of existing requests. If all requests have long past tokens, then the capacity is likely to be low (See Figure˜2). Then there is no way to define a fixed amount of equal share.
@@
The second challenge is the characteristic of unknown output length before finishing a request. This prevents a direct adaptation of classical algorithms like SFQ and Deficit Round Robin (DRR) [41] into the LLM serving. SFQ-style algorithms can provide good bounds in fairness by setting the Start and Finish tags through virtual time, as introduced in Section˜2.2. However, computing Start and Finish tags requires knowing the request length in advance. DRR performs round-robin scheduling with a “deficit counter” mechanism to achieve fair scheduling of packets of variable length. In DRR, each client is assigned a specific quantum of service. It tracks the “deficit” of service for each client to ensure fairness over time. During each round, the scheduler allows each client to dispatch as many requests as possible, provided that the total length of these requests does not exceed the sum of the client’s assigned quantum for that round and any accumulated deficit from previous rounds. Without knowing the length in advance, DRR cannot determine how many jobs can be scheduled within the quantum.
We will give the definition for LLM serving fairness in Section˜3, and give an algorithm to achieve the LLM serving fairness in Section˜4.
@@ 3 Definition of Fairness in LLM Serving
As discussed in Section˜2.3, how to define the cost of a request in LLM serving is not obvious. In this section, we discuss how to define the cost for a request, and how to measure the service a client has received (Section˜3.1). After defining the measurement of service, we can define fairness among clients in Section˜3.2.
@@ 3.1 Measurement of Service
In this subsection, we will discuss the measurement of the service a client has received. Specifically, we define $W_{f}(t_{1},t_{2})$ and $W_{g}(t_{1},t_{2})$ from Equation˜1 in the context of LLM serving. We omit the subscript, $W(t_{1},t_{2})$, when the client is clear from the context or irrelevant.
@@
[…]
However, both of these formulations cannot accurately reflect the actual LLM serving cost: The computation of the tokens at the prefill stage can be parallelized and achieve high GPU utilization. However, at the generation stage, we can only generate tokens one by one, as each token depends on all previous tokens as described in Section˜2.1.
Weighted number of tokens To better reflect the actual LLM serving cost, a more accurate measure should capture the difference in costs of the prefilling and generation phases. One simple way to implement this idea is by using a weighted combination of the prefilling (input) tokens and decoding (output) tokens, inspired by the pricing mechanism used in OpenAI’s API [^3] https://openai.com/pricing. Formally, let $w_{p}$ be the weight of input tokens and $w_{q}$ be the weight of output tokens. Then, we have $W(t_{1},t_{2})=w_{p}\cdot n_{p}+w_{q}\cdot n_{q}$, where $n_{p},n_{q}$ are the number of processed input and output tokens, respectively. Following OpenAI pricing, we set $w_{p}=1$ and $w_{q}=2$. Due to its simplicity, we will use this measure in our analysis and evaluation.
Customized, unified representation. The definition of fairness in LLM serving can also be extended to other aspects, such as the weighted number of FLOPs or a more sophisticated method introduced in [27] that uses piecewise linear functions for the number of input and output tokens. The important point to note here is that we need to separate the cost for input and output tokens. Generally, the service can be represented as a function of the number of input and output tokens ($l_{int},l_{out}$, respectively). Let $h(l_{in},l_{out})$ be the cost function that is monotonically increasing according to $l_{in}$ and $l_{out}$. Our method can easily accommodate to different $h$ (Section˜4.2).
@@ 3.2 Fairness in LLM Serving
The first property means that two clients sending requests at more than their fair share will get the same service, regardless of the discrepancy between their sending rates. The second property says that a client sending requests at a higher service rate will not get less service than a client sending at a lower service rate. Basically, the first two properties say that a misbehaving client is contained (i.e., doesn’t receive more service than other backlogged clients), not punished (i.e., doesn’t receive less service than other non-backlogged clients). Finally, the work conserving property aims to maximize the utilization, addressing a key weakness of the RPM-based solutions.
The three properties above assume an idealized fair serving system. A practical system will invariably approximate these properties. In general, the best we can achieve is deriving bounds that are independent of the length of the time interval, e.g., in the first property, the difference between $W_{f}(t_{1},t_{2})$ and $W_{g}(t_{1},t_{2})$ is bounded by a value that is independent of $t_{2}-t_{1}$. We give the formal guarantees provided by our method in Section˜4.1.

### v2

@@ Abstract
High-demand LLM inference services (e.g., ChatGPT and BARD) support a wide range of requests from short chat conversations to long document reading. To ensure that all client requests are processed fairly, most major LLM inference services have request rate limits, to ensure that no client can dominate the request queue. However, this rudimentary notion of fairness also results in under-utilization of the resources and poor client experience when there is spare capacity. While there is a rich literature on fair scheduling, serving LLMs presents new challenges due to their unpredictable request lengths and their unique batching characteristics on parallel accelerators. This paper introduces the definition of LLM serving fairness based on a cost function that accounts for the number of input and output tokens processed. To achieve fairness in serving, we propose a novel scheduling algorithm, the Virtual Token Counter (VTC), a fair scheduler based on the continuous batching mechanism. We prove a 2×2$\times$2 × tight upper bound on the service difference between two backlogged clients, adhering to the requirement of work-conserving. Through extensive experiments, we demonstrate the superior performance of VTC in ensuring fairness, especially in contrast to other baseline methods, which exhibit shortcomings under various conditions. The reproducible code is available at https://github.com/Ying1123/VTC-artifact.
@@ 1 Introduction
Today’s LLM serving systems [24, 20] typically use First-Come-First-Serve (FCFS) to schedule incoming requests. While simple, this scheduling discipline has several drawbacks. One such drawback is the lack of isolation: a client sending a disproportionate number of requests can negatively impact the service of all the other clients sharing the same server (i.e., slow down their requests or even cause timeouts) even when they send very little traffic. In multi-tenant personalized serving (S-LoRA [43], Punica [8]) that uses a dedicated adapter for each user, it is important to ensure fairness among the adapters as well. One solution to address this problem is to limit the incoming load of each client. Many of the existing LLM services do this today by imposing a request-per-minute (RPM) limit [37] for each client.
@@
This is a common problem in many other domains like networking and operating systems. The solution of choice to achieve both isolation and high resource utilization in those domains has been fair queueing [30]. Fair queueing ensures that each client will get their “fair share”. In the simplest case, if there are $n$ clients sharing the same resource, the fair share is at least $1/n$ of the resource, which means that each client gets at least $1/n$ of the resource. Furthermore, if some clients do not use their share, other clients with more demands can use it, hence leading to higher resource utilization.
@@
Despite these similarities, we cannot directly use the algorithms developed for networking and operating systems, as LLM serving has several unique characteristics. First, the request output lengths are unknown in advance. In contrast, in networking, the packet lengths are known before the packet is scheduled. Second, the cost of each token can vary. For instance, the cost of processing an input (prompt) token is typically lower than that of an output token, because input token processing is parallelizable. In contrast, the cost of sending a bit or the cost of a CPU time slice are the same irrespective of the workload. Third, the effective capacity of an LLM server (i.e., processing rate expressed in token/sec when the request queue is non-empty) can vary over time. For example, longer input sequences take more memory. This limits the number of batched parallel requests during generation, leading to GPU under-utilization and a lower processing rate. In contrast, the network or CPU capacity is assumed to be fixed.
In this paper, we discuss the factors that need to be considered when defining fairness in the context of LLM serving. We show how different definitions can be incorporated into a configurable service cost function in Section 3. While the cost function can be customized, a simple metric of counting input and output tokens at different prices is extensively used in analysis for the sake of simplicity. We then present a fair scheduling algorithm called Virtual Token Counter (VTC) that can be easily adapted for different service cost functions. At a high level, VTC tracks the services received for each client and will prioritize the ones with the least services received, with a counter lift each time a client is new to the queue. It updates the counters at a token-level granularity on the fly, which addresses the unknown length issue. VTC integrates seamlessly with current LLM serving batching techniques (Section 2.1), and its scheduling mechanism does not depend on the server’s capacity, overcoming the problem of the dynamically fluctuating server capacity. We also provide theoretical bounds of fairness for VTC in Section 4.1. The serving architecture of VTC is illustrated in Figure 1.
@@
  • This is the first work to discuss the fair serving of Large Language Models to the best of our knowledge. We identify its unique challenges and give the definition of LLM serving fairness (Section 3).
  • We propose a simple yet effective fair-serving algorithm called VTC. We provide rigorous proofs for VTC on fairness guarantee, which gives fairness bound within 2×2$\times$2 × of the optimal bound (Section 4).
  • We conduct in-depth evaluations on our proposed algorithm VTC. Results confirm that our proposed algorithms are fair and work-conserving (Section 5).
@@ 2.1 Large Language Models Serving
The generation procedure consists of two stages: the initial prefilling stage, and the decoding stage [39]. Mathematically, $x$ is a sequence of tokens $(x_{1},x_{2},...,x_{n})$. In the prefilling stage, the LLM computes the probability of the first new tokens: $P(x_{n+1}|x_{1},...,x_{n})$. In the decoding stage, the system autoregressively generates a new token. At time $t$ ($t\geq 1$), the process is written as: $P(x_{n+t+1}|x_{1},...,x_{n+t})$.
@@
Figure 1: Serving architecture with Virtual Token Counter (VTC), illustrated with two clients. VTC maintains a queue of requests and keeps track of tokens served for each client. In each iteration of the LLM execution engine, some tokens from some clients are generated. The counters of these clients are correspondingly updated. When the condition of adding new requests is satisfied (e.g. memory is released when some other requests finish), VTC will be invoked to choose the requests to be added. VTC achieves fairness by prioritizing clients with the lowest counter and carefully handling clients’ leave and rejoin (Section 4.1).
@@
Naively, the execution stream can choose to execute requests one by one. However, this is highly GPU inefficient due to various natures of the LLM generation procedure. For instance, the decoding steps must be carried out sequentially where the arithmetic intensity is relatively low in a single step. Contemporary serving systems usually perform batching that executes multiple requests concurrently to maximize the system throughput. The most widely used approach in LLM serving is continuous batching [50]. Algorithm 1 shows the pseudocode for continuous batching. [^2] For a simple presentation, we consider an implementation that only uses continuous batching for decode steps but keeps the prefill step separated. as how TGI [21] adopted the original proposed iteration-level scheduling in Orca [50]. For more discussions, see Section C.1. The monitoring stream enqueues requests to a waiting queue. The execution stream performs a check on whether there are finished requests at the end of each decoding step. If there are, the system removes these requests and adds new requests from the queue.
@@
We adopt a continuous batching scheme in which a request only leaves the batch when it generates an EOS token or reaches the pre-defined maximum number of generated tokens (i.e., no preemption). This paper focuses on integrating fair scheduling with continuous batching, and we leave an investigation on preemption as an orthogonal future work (discussed in Section C.3).
@@ 2.2 Existing Fairness Approaches
All fair queueing algorithms maintain a system virtual time, $v(t)$, which intuitively measures the service received by a continuously backlogged flow in terms of bits forwarded. Each packet, $p$ is associated two tags: Start tag $S(p)$ and a Finish tag $F(p)=S(p)+l_{p}$. The Start tag (a.k.a. packet’s virtual starting time) is computed based on both the system virtual time and the Finish tag (a.k.a. packet’s virtual finishing time). These algorithms schedule packets in the ascending order of either the Finish or Start tags.
@@
There exists a distinct strand of research [6, 3, 47] focusing on the fair scheduling of preemptible tasks (e.g., CPU scheduling). The Completely Fair Scheduler (CFS) [1], implemented in Linux 2.6.23 and applying fair queuing to CPU scheduling, is closely related to our algorithm. In CFS, a “vruntime” is maintained for each task, and the task with the smallest “vruntime” is scheduled next. The tasks can be presented with a small time slice, aiming to maximize overall CPU utilization while also maximizing interactive performance.
@@ 2.3 Challenges
Additionally, in LLM serving, the server has variable token-rate capacity, although the memory allocated for a batch is constant. Firstly, even if the request queue is not empty, we are not guaranteed that each batch is full. This is because we need to preserve spaces for future generated tokens, and also because the tokens added to the batch are not at the token but the request granularity. Secondly, the number of tokens processed highly depends on the requests’ arrival patterns because of the continuous batching mechanism (Section 2.1). Furthermore, the capacity depends on the mix between input and output tokens of existing requests. If all requests have long past tokens, then the capacity is likely to be low (See Figure 2). Then there is no way to define a fixed amount of equal share.
@@
The second challenge is the characteristic of unknown output length before finishing a request. This prevents a direct adaptation of classical algorithms like SFQ and Deficit Round Robin (DRR) [45] into the LLM serving. SFQ-style algorithms can provide good bounds in fairness by setting the Start and Finish tags through virtual time, as introduced in Section 2.2. However, computing Start and Finish tags requires knowing the request length in advance. DRR performs round-robin scheduling with a “deficit counter” mechanism to achieve fair scheduling of packets of variable length. In DRR, each client is assigned a specific quantum of service. It tracks the “deficit” of service for each client to ensure fairness over time. During each round, the scheduler allows each client to dispatch as many requests as possible, provided that the total length of these requests does not exceed the sum of the client’s assigned quantum for that round and any accumulated deficit from previous rounds. Without knowing the length in advance, DRR cannot determine how many jobs can be scheduled within the quantum. Compared to CPU scheduling, although exploring adequate preemption is worthwhile in LLM serving, it cannot occur frequently. We need to define service fairness in LLM serving and operate at the granularity of individual tokens when frequent preemption is not possible. Additionally, the Completely Fair Scheduler (CFS) in CPU scheduling does not account for the concurrency of each task. It seeks fairness among individual tasks rather than among streams of tasks that can be executed concurrently.  [^3] Our algorithm does not consider preemption. Discussion about preemption is in Section C.3.
We will give the definition for LLM serving fairness in Section 3 and give a scheduling algorithm to achieve the LLM serving fairness in Section 4. We will outline our algorithm in a basic format for clarity, while details on its general form and integration with existing serving frameworks can be found in Section C.1. Although our algorithm is closely related to CFS, we also discuss the adaptation of DRR in Section C.2. Further discussions on future work are included in Section C.3.
@@ 3 Definition of Fairness in LLM Serving
In this section, we discuss the cost of a request, and the measurement of the service a client has received (Section 3.1). After defining the measurement of service, we can define fairness among clients in Section 3.2.
Notation Explanation
$W_{f}(t_{1},t_{2})$ service received by $f$ during interval $[t_{1},t_{2})$ (write as $W(t_{1},t_{2})$ when $f$ is clear in the context)
$n_{p}$ number of processed input tokens
$n_{q}$ number of processed output tokens
$w_{p}$ weight of input tokens in the cost function
$w_{q}$ weight of output tokens in the cost function
$h(n_{p},n_{q})$ customized cost function
$c_{i}$ virtual token counter for client $i$
$Q$ waiting queue of requests to be processed
$i\in Q$ $\exists r\in Q$, $r$ is a request from client $i$
$L_{input}$ maximum number of input tokens in a request
$L_{output}$ maximum number of output tokens in a request
$M$ maximum number of tokens that can be fitted in a running batch
$U$ invariant bound: $\max(w_{p}\cdot L_{input},w_{q}\cdot M)$
Table 1: The upper half includes notations for service measurement. The lower half includes notations for the VTC algorithm and its analysis. The terms $n_{p},n_{q}$ can refer to either a single request or a single client, depending on the context.
@@ 3.1 Measurement of Service
In this subsection, we discuss the measurement of the service a client has received. Specifically, we define $W_{f}(t_{1},t_{2})$ and $W_{g}(t_{1},t_{2})$ from Equation 1 in the context of LLM serving. We omit the subscript and write $W(t_{1},t_{2})$ when the client is clear from the context or is irrelevant. The number of processed input and output tokens are denoted as $n_{p},n_{q}$. Notations that will be introduced and used multiple times in this paper are summarized in Table 1.
Number of tokens A straightforward way to measure the service provided to a client is by summing the number of input tokens that have been processed and the number of output tokens that have been generated so far, i.e., $W(t_{1},t_{2})=n_{p}(t_{1},t_{2})+n_{q}(t_{1},t_{2})$ during the time window $[t_{1},t_{2})$.
@@
However, both of these formulations cannot accurately reflect the actual LLM serving cost: The computation of the tokens at the prefill stage can be parallelized and achieve high GPU utilization. However, at the generation stage, we can only generate tokens one by one, as each token depends on all previous tokens as described in Section 2.1.
Weighted number of tokens To better reflect the actual LLM serving cost, a more accurate measure should capture the difference in costs of the prefilling and generation phases. One simple way to implement this idea is by using a weighted combination of the prefilling (input) tokens and decoding (output) tokens, inspired by the pricing mechanism used in OpenAI’s API [^4] https://openai.com/pricing. Formally, let $w_{p}$ be the weight of input tokens and $w_{q}$ be the weight of output tokens. Then, we have $W(t_{1},t_{2})=w_{p}\cdot n_{p}(t_{1},t_{2})+w_{q}\cdot n_{q}(t_{1},t_{2})$. Due to its simplicity, we will use this measure extensively in our analysis and evaluation.
Customized, unified representation. The definition of fairness in LLM serving can also be extended to other aspects, such as the weighted number of FLOPs or a more sophisticated method introduced in [31] that uses piecewise linear functions for the number of input and output tokens. Generally, the service can be represented as a function of the number of input and output tokens ($n_{p},n_{q}$, respectively). Let $h(n_{p},n_{q})$ be the cost function that is monotonically increasing according to $n_{p}$ and $n_{q}$. Our method can easily accommodate different $h$ (Section 4.2).
@@ 3.2 Fairness in LLM Serving
The first property means that two clients sending requests at more than their fair share will get the same service, regardless of the discrepancy between their sending rates. The second property says that a client sending requests at a higher rate will not get less service than a client sending at a lower rate. Basically, the first two properties say that a misbehaving client is contained (i.e., doesn’t receive more service than other backlogged clients), and not punished (i.e., doesn’t receive less service than other non-backlogged clients). Finally, the work conserving property aims to maximize the utilization, addressing a key weakness of the RPM-based solutions.
The three properties above assume an idealized fair serving system. A practical system will approximate these properties. In general, the best we can achieve is deriving bounds that are independent of the length of the time interval, e.g., in the first property, the difference between $W_{f}(t_{1},t_{2})$ and $W_{g}(t_{1},t_{2})$ is bounded by a value that is independent of $t_{2}-t_{1}$. We give the formal guarantees provided by our method in Section 4.1.